# Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong Lens

### v1

@@ Appendix A Optimal Hyperparameter Values for the Matern Kernel
$(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ $4.2074^{+0.1840}_{-0.2042}$ $-0.0091^{+0.0805}_{-0.0542}$ $6.1800^{+2.4526}_{-2.0224}$
@@
Table 4: Optimal hyperparameter values for regularising lensing potential perturbations using the Matérn kernel, determined objectively by sampling the Bayesian evidence. Results are shown for inversions of the lensing potential only ($\delta\boldsymbol{\psi}$-only), joint inversions of the source and lensing potential using the true main lens and source models ($(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$), and joint inversions starting from the best-fit macro model without potential perturbations ($(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$). Superscripts and subscripts represent the $68\%$ credible interval.
Appendix B Robustness Analysis of the Potential Correction Results for the Jackpot Lens
@@
Figure 14: Convergence correction results for two Jackpot-like mock lenses, demonstrating the ability of our potential correction method to reveal the mass distribution of the input subhalos. The left lens includes an input NFW subhalo with a mass of $\sim 10^{10}\,M_{\odot}$ and a high concentration ($\sim 200$), centred at $(-1.040,-0.651)$ (Despali et al., 2024). The right lens contains an input NFW subhalo with a mass of $\sim 10^{11.5}\,M_{\odot}$, following the standard mass-concentration relation, centred at $(-1.00,-1.10)$ (Nightingale et al., 2024).
@@
In addition to the localised mass clumps corresponding to the subhalo, additional mass features are observed in the bottom-right panel of Figures 10 and  12, including a dipole-like structure near the detected subhalo and a negative convergence peak at $(x,y)=(-1.0,-1.0)$. To investigate whether these additional mass features are genuine or induced by systematic errors in our potential correction method, we adopt a parametric approach analogous to that employed by Despali et al. (2024) in modelling the Jackpot lens. This approach incorporates an EPL plus shear model for the mass distribution of the main lens, an NFW model for the subhalo, and a pixelized source model. Subsequently, we simulate 1000 Jackpot-like mock lenses based on this model, each with a distinct realisation of image noise. Applying our potential correction method to these mock lenses consistently reproduces the additional mass features observed in the analysis of the real Jackpot lens data. We then utilise the median and standard deviation computed from the 1000 convergence perturbation maps (derived from the 1000 Jackpot-like mock lenses) as proxies for the signal and noise of the reconstructed mass perturbation features. The ratio between these quantities defines the signal-to-noise ratio (SNR), which quantifies the statistical significance of the reconstructed mass perturbation features. We find that the SNR of these additional mass features exceeds 3, confirming that they represent genuine artefacts induced by a systematic error in our potential correction method rather than being introduced by random noise (noting that these additional mass features should be absent in the mock lens data). If the potential correction is initiated using the true main lens mass model and source light model, the additional mass features vanish from the correction result, as illustrated in the bottom-right panel of Figure 13. We thus conclude that these additional features result from initiating the potential correction with a biased estimate of the main lens mass and source light provided by the macro model. Notably, the convergence correction shown in the bottom-right panel of Figure 13 recovers the central high-density region of the cuspy subhalo more accurately than the correction initiated from a biased estimate (see Figure 11). Hence, in addition to the limited dynamical range arising from the use of constant regularisation in the pixelized potential correction (discussed in Section 5.1), initiating the correction with a biased main lens mass and source estimation also contributes to the underestimation of the central convergence of a cuspy subhalo.

### v2

@@ Appendix B Optimal Hyperparameter Values for the Matern Kernel
$(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$ $4.2074^{+0.1840}_{-0.2042}$ $0.0091^{+0.0805}_{-0.0542}$ $6.1800^{+2.4526}_{-2.0224}$
@@
Table 4: Optimal hyperparameter values for regularising lensing potential perturbations using the Matérn kernel, determined objectively by sampling the Bayesian evidence. Results are shown for inversions of the lensing potential only ($\delta\boldsymbol{\psi}$-only), joint inversions of the source and lensing potential using the true main lens and source models ($(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Truth}}$), and joint inversions initialised from the point estimate of the macro model without potential perturbations ($(\delta\boldsymbol{\psi},\boldsymbol{s})^{\mathrm{Macro}}$). Superscripts and subscripts represent the $68\%$ credible interval.
Appendix C Robustness Analysis of the Potential Correction Results for the Jackpot Lens
@@
Figure 14: Convergence correction results for two Jackpot-like mock lenses, demonstrating the ability of our potential correction method to reveal the mass distribution of the input subhalos. The left lens includes an input NFW subhalo with a mass of $\sim 10^{10}\,M_{\odot}$ and a high concentration ($\sim 200$), centred at (1.040,0.651)(-1.040,-0.651) (Despali et al., 2024). The right lens contains an input NFW subhalo with a mass of $\sim 10^{11.5}\,M_{\odot}$, following the standard mass-concentration relation, centred at (1.00,1.10)(-1.00,-1.10) (Nightingale et al., 2024).
@@
In addition to the localised mass clumps corresponding to the subhalo, additional mass features are observed in the bottom-right panel of Figures 10 and  12, including a dipole-like structure near the detected subhalo and a negative convergence peak at $(x,y)\approx(1.0,1.0)$. To investigate whether these additional mass features are genuine or induced by systematic errors in our potential correction method, we adopt a parametric approach analogous to that employed by Despali et al. (2024) in modelling the Jackpot lens. This approach incorporates an EPL plus shear model for the mass distribution of the main lens, an NFW model for the subhalo, and a pixelized source model. Subsequently, we simulate 1000 Jackpot-like mock lenses based on this model, each with a distinct realisation of image noise. Applying our potential correction method to these mock lenses consistently reproduces the additional mass features observed in the analysis of the real Jackpot lens data. We then utilise the median and standard deviation computed from the 1000 convergence perturbation maps (derived from the 1000 Jackpot-like mock lenses) as proxies for the signal and noise of the reconstructed mass perturbation features. The ratio between these quantities defines the signal-to-noise ratio (SNR), which quantifies the statistical significance of the reconstructed mass perturbation features. We find that the SNR of these additional mass features exceeds 3, confirming that they represent genuine artefacts induced by a systematic error in our potential correction method rather than being introduced by random noise (noting that these additional mass features should be absent in the mock lens data). If the potential correction is initiated using the true main lens mass model and source light model, the additional mass features vanish from the correction result, as illustrated in the bottom-right panel of Figure 13. We thus conclude that these additional features result from initiating the potential correction with a biased estimate of the main lens mass and source light provided by the macro model. Notably, the convergence correction shown in the bottom-right panel of Figure 13 recovers the central high-density region of the cuspy subhalo more accurately than the correction initiated from a biased estimate (see Figure 11). Hence, in addition to the limited dynamical range arising from the use of constant regularisation in the pixelized potential correction (discussed in Section 5.1), initiating the correction with a biased main lens mass and source estimation also contributes to the underestimation of the central convergence of a cuspy subhalo.